FGCONSOLE(1) LinuxUser’sManual FGCONSOLE(1) NAME fgconsole−printthenumberoftheactiveVT. SYNOPSIS fgconsole DESCRIPTION IftheactiveVTis/dev/ttyN,thenprintsNonstandardoutput. SEEALSO chvt(1). Consoletools 15Oct1997 1